# The Impact of VR and 2D Interfaces on Human Feedback in Preference-Based Robot Learning

### v1

@@ Abstract
Aligning robot navigation with human preferences is essential for ensuring comfortable and predictable robot movement in shared spaces, facilitating seamless human-robot coexistence. While preference-based learning methods, such as reinforcement learning from human feedback (RLHF), enable this alignment, the choice of the preference collection interface may influence the process. Traditional 2D interfaces provide structured views but lack spatial depth, whereas immersive VR offers richer perception, potentially affecting preference articulation. This study systematically examines how the interface modality impacts human preference collection and navigation policy alignment. We introduce a novel dataset of 2,325 human preference queries collected through both VR and 2D interfaces, revealing significant differences in user experience, preference consistency, and policy outcomes. Our findings highlight the trade-offs between immersion, perception, and preference reliability, emphasizing the importance of interface selection in preference-based robot learning. The dataset will be publicly released to support future research.
@@ I Introduction
Despite the potential advantages of VR, systematic comparisons between immersive and traditional 2D video-based interfaces for preference collection remain sparse. Previous studies have primarily focused on individual interface performance or user engagement without thoroughly investigating how the chosen interface modality affects preference data quality and the subsequent alignment of robot navigation policies. A critical question arises: Do user preferences for robot behavior differ between immersive VR and traditional 2D interfaces, and if so, how do these differences influence preference-aligned navigation policies?
Figure 1: Left: Our study collects preferences via Virtual Reality (VR) and 2D interfaces (top-down & first-person views), enabling a systematic comparison of interface modalities for eliciting human preferences on robot navigation. Top-right: The preference dataset consists of 2,325 navigation queries from 31 participants. Bottom-right: Using Reinforcement Learning from Human Feedback (RLHF) with our dataset, we refine a standard navigation policy (blue) into preference-aligned policies (orange), enhancing human-centric navigation.
In this paper, we introduce a novel user preference dataset for robot navigation, collected through both VR and 2D interfaces. Our study systematically evaluates the impact of interface modality on the user experience and preference collection process. We compare preferences elicited using a VR interface against those obtained via first-person and bird’s-eye/top-down view video interfaces. Finally, we derive preference-aligned navigation policies from the collected preference data, taking into account the interface modality used for collection. Our primary contributions are:
@@
  • A quantitative and qualitative evaluation of how the different interface modalities influence the collection of user preferences.
  • The derivation and analysis of three preference-aligned navigation policies from modality-specific user preferences.
@@ III Method
We subsequently provide an overview about the navigation task of the robot, the interface, and user study setup for data collection.
@@ III-B Learning Robot Navigation
In line with our target methodology, PbRL, the navigation task is solved via reinforcement learning, where an agent learns to navigate the robot to the goal using velocity commands. As a simulation environment, iGibson [28] with its PyBullet physics engine is used to simulate a Kobuki Turtlebot 2i. The navigation scene resembles an open space with position-randomized small and large box obstacles and a static human. Start, goal and human position are sampled in in close proximity to each other. The core navigation reward contains a sparse goal reward, a continuous time penalty, sparse collision and timeout penalties, and penalties for self-intersecting trajectories and jerk:
@@ III-C Query Generation
Once trained, the ensemble $\mathcal{E}$ is used to generate diverse trajectory options for the randomized scene configuration, by sampling two individual ensemble policies and rolling them out in the sampled navigation scenario. To ensure meaningful queries, trajectories that result in collisions or self-intersections are filtered out. We generate a dataset $D_{Q}$ of $N=500$ queries, from which we sample subsets for the participants to rate in the user study, either in VR or as 2D video playback.
@@ III-D Query Interfaces
To collect user preferences, we employ three types of query interfaces: an immersive virtual reality (VR) setup and two 2D video-based KBM interfaces on a desktop computer. Both interfaces show the same navigation environment but differ in perspective and immersion. The VR interface provides an interactive and immersive experience, whereas the 2D video KBM interface offers a more conventional, screen-based alternative. Within the 2D interface, we present two distinct perspectives: Top-down (2D-TD), as in [24], also known as a bird’s-eye view, and first-person view (2D-FPV), which more closely resembles a VR perspective. For a given query, the human position is defined, serving additionally as the observer position in VR and 2D-FPV.
@@ III-D1 Virtual Reality
[…]
@@
[…]
@@ III-E User Study
We conducted a user study to compile a dataset of participants’ navigation preferences and assess user experience across three preference interfaces (VR, 2D-TD, 2D-FPV). Data collection was divided into three distinct stages: S1: Collecting preferences through navigation queries for each interface modality, S2: Post-interface interaction questionnaires assessing user experience, and S3: A final ranking survey comparing the three interfaces. Before testing, all participants received detailed study information, provided written consent, and completed a demographic questionnaire. The study was structured into three blocks in randomized order, each corresponding to one of the interfaces as an experimental condition. Each block presented the same 25 preference queries in random order (S1), initially sampled for each participant from the query dataset $D_{Q}$. By presenting identical queries across different interfaces, we could later investigate the impact of the interface modality on participants’ navigation preferences. After each interface block, participants completed a questionnaire (S2) regarding their experience with the interface and the queries, as shown in I, Questions Q1–Q10. Upon completion of all three blocks, the final ranking survey (S3), based on the Technology Acceptance Model (TAM) [32], asked participants to rank the interfaces by perceived usefulness, intention to use, and ease of use, as shown in I, Questions R1–R3. All collected data were anonymized using a coding table for participant IDs. Each session lasted approximately one hour.
@@ IV Experimental Evaluation
We propose the following hypotheses that we aim to evaluate with our study: (H1) The user experience differs between the interface modalities. (H2) The user preferences for robot navigation differ between the interface modalities. (H3) A preference discrepancy between interfaces reflects in the navigation behavior of interface-specific preference-aligned policies.
@@ IV-A Interface Questionnaire
[…]
@@ IV-C Dataset Overview
The dataset contains 2,325 user preference queries for robot navigation, collected across three interface modalities: VR, 2D-TD, and 2D-FPV. The dataset’s structure follows the data collection design: Participant $\rightarrow$ interface modality $\rightarrow$ query $\rightarrow$ preference label.
In more detail, each query stores both trajectories and the scene configuration (e.g., for replay in VR), the 2D videos, and RL episodes (state, action, next state, reward) for both trajectories, where states $\mathbf{s}_{t}=(\text{lidar},\text{goal},\text{human})$ capture lidar data, robot-centric goal and human position, while actions $\mathbf{a}_{t}=(v,\omega)$ represent velocity commands. Preferences are stored as A/B labels. In addition to the state-action pairs, we provide the robot trajectory as a 2D path, the static human pose, and the obstacle poses in world coordinates.
As we later show, our dataset enables the distillation of preference models, e.g., for the preference alignment of RL robot navigation policies. The dataset will be published upon paper acceptance.
@@ IV-D2 Disagreement Analysis
In cases where participants exhibited different preferences for the same queries in VR and on 2D interfaces, participants preferred shorter and more straightforward driving trajectories (measured by the area under the path) in 2D compared to VR. Additionally, the robot may traverse closer to the human, compare (e). Similarly, query disagreements point to a preference for more straightforward driving styles when transitioning from the 2D-FPV to the 2D-TD interface, see (f).
@@ IV-E Policy Alignment
Steps 49.1 49.5 55.9 50.0
Path Length [m] 4.9 4.9 5.4 4.9
Time [s] 9.6 9.7 11.0 9.8
AA [rad] 3.2 3.1 4.1 2.8
AUP [m²] 2.2 1.8 3.4 2.1
$\min(d_{h})$ [m] 1.33 1.20 1.28 1.25
Speed $\bar{v}$ [m/s] 0.51 0.50 0.49 0.51
TABLE III: Quantitative metrics for our preference-aligned $\pi_{\text{VR}}$, $\pi_{\text{\mbox{2D-TD}}}$, and $\pi_{\text{\mbox{2D-FPV}}}$ and a non-aligned baseline $\pi_{\text{BL}}$, averaged over 1,000 trajectories in randomly sampled scene configurations.
To examine how preferences collected from different interfaces impact policy alignment and navigation behavior, we employ a preference-based reinforcement learning (PbRL) approach. This method trains reward functions from the collected preference dataset, enabling the subsequent alignment of human-aware RL navigation policies for each query interface. Following [10, 9], we learn a parametric reward function $\hat{r}_{\psi}$ from human preferences. For policy optimization, we implement a PbRL algorithm following [10], using TD3 [29] as the base RL algorithm. Splitting up our preference dataset by interface, we train three individual reward functions modeled as an MLP with $[256,256,256]$ hidden units based on the queries collected from the participants. The policies for each condition were then trained for 500k time steps by weighting the original navigation task reward and their respective preference reward model as $r=\lambda\hat{r}_{\psi}+(1-\lambda)r_{\text{core}}$, with $\lambda=0.2$. The value of $\lambda$ was empirically determined through iterative experimentation to ensure training stability. Note that the preference reward model ranges from -1 to +1, while the original task reward ranges from a minimum of -10 to a maximum of +20, with these extrema driven by the sparse terms in the function. In addition to the trained policies $\pi_{\text{VR}}$, $\pi_{\text{\mbox{2D-TD}}}$, and $\pi_{\text{\mbox{2D-FPV}}}$, we also include a non-aligned baseline policy $\pi_{\text{BL}}$, trained on the same task without preference-based rewards, for comparison.
@@
The quantitative results in Table III confirm that preference-aligned policies improve human-aware navigation, with $\pi_{\text{VR}}$ achieving the best balance between efficiency and safety by maintaining the highest human clearance (1.33 m) while ensuring low travel times. $\pi_{\text{\mbox{2D-FPV}}}$ prioritizes safety, exhibiting the lowest collision rate (4.9%) but at the cost of longer paths (5.4 m) and higher angular accumulation (4.1 rad). $\pi_{\text{\mbox{2D-TD}}}$ exhibits the most straight-forward navigation, reflected in the lowest area under path (AUP). These findings, supported by the qualitative analysis, highlight that $\pi_{\text{VR}}$ offers the most balanced navigation performance, while $\pi_{\text{\mbox{2D-FPV}}}$ navigates overly conservatively, and $\pi_{\text{BL}}$ is less cautious. Consequently, we find empirical support for H3.
@@ V Conclusion
Our study systematically examined how interface modality affects human preference collection for robot navigation and the resulting policy alignment. The results confirm that the choice of interface modality significantly impacts how users express preferences, perceive the interaction, and ultimately shape robot navigation behavior. The VR interface provided a more immersive and intuitive experience, leading to greater confidence and ease in preference expression. This aligns with prior findings on the benefits of immersive environments for user engagement [35]. However, the study also revealed that preferences were not entirely consistent between interfaces, with participants favoring shorter, more direct paths in 2D interfaces while exhibiting greater tolerance for curved trajectories with increased human clearance in VR. This suggests that the visualization and spatial representation of the robot’s movement influence user preferences, highlighting the importance of maintaining interface consistency during preference collection. The navigation policies trained on interface-specific preferences demonstrated noticeable differences in robot behavior. These findings highlight the necessity of considering interface effects in preference-based reinforcement learning, as user preference shifts due to modality changes can directly impact policy training outcomes. To support further research on these effects and the collected user preferences, we will make the preference dataset available to the research community upon paper acceptance.

### v2

@@ Abstract
Aligning robot navigation with human preferences is essential for ensuring comfortable, and predictable robot movement in shared spaces. While preference-based learning methods, such as reinforcement learning from human feedback (RLHF), enable this alignment, the choice of the preference collection interface may influence the process. Traditional 2D interfaces provide structured views but lack spatial depth, whereas immersive VR offers richer perception, potentially affecting preference articulation. This study systematically examines how the interface modality impacts human preference collection and navigation policy alignment. We introduce a novel dataset of 2,325 human preference queries collected through both VR and 2D interfaces, revealing significant differences in user experience, preference consistency, and policy outcomes. Our findings highlight the trade-offs between immersion, perception, and preference reliability, emphasizing the importance of interface selection in preference-based robot learning. The dataset is available to support future research.
@@ I Introduction
Despite the potential advantages of VR, systematic comparisons between immersive and traditional 2D video-based interfaces for preference collection remain sparse. Previous studies have primarily focused on individual interface performance or user engagement without thoroughly investigating how the chosen interface modality affects preference data quality and the subsequent alignment of robot navigation policies. A critical question is whether user preferences differ between VR and 2D interfaces, and how these differences influence navigation policies.
Figure 1: Left: Our study collects preferences via Virtual Reality (VR) and 2D interfaces (top-down & first-person views), enabling a systematic comparison of interface modalities for collecting human preferences on robot navigation. Top-right: The preference dataset consists of 2,325 navigation queries from 31 participants. Bottom-right: Using Reinforcement Learning from Human Feedback (RLHF) with our dataset, we refine a standard navigation policy (blue) into preference-aligned policies (orange).
In this paper, we introduce a novel user preference dataset for robot navigation, collected through both VR and 2D interfaces, see Fig. 1. Our study systematically evaluates the impact of interface modality on the user experience and preference collection process. We compare preferences elicited using a VR interface against those obtained via first-person and bird’s-eye/top-down view video interfaces. Finally, we derive preference-aligned navigation policies from the collected preference data, taking into account the interface modality used for collection. Our primary contributions are:
@@
  • A quantitative and qualitative evaluation of how different interfaces influence the collection of user preferences.
  • The derivation and analysis of three preference-aligned navigation policies from modality-specific preferences.
@@ III Method
We subsequently provide an overview on the robot’s navigation task, the interface, and user study for data collection.
@@ III-B Learning Robot Navigation
In line with our target methodology, PbRL, the navigation task is solved via reinforcement learning, where an agent learns to navigate the robot to the goal using velocity commands. As a simulation environment, iGibson [28] with its PyBullet physics engine is used to simulate a Kobuki Turtlebot 2i. The navigation scene resembles an open space with position-randomized small and large box obstacles and a static human. Start, goal and human position are sampled in in close proximity to each other. The core navigation reward contains a sparse goal reaching reward, a continuous time penalty, sparse collision and timeout penalties, and penalties for self-intersecting trajectories and jerk:
@@ III-C Query Generation
Once trained, the ensemble $\mathcal{E}$ is used to generate diverse trajectory options for the randomized scene configuration, by sampling two individual ensemble policies and rolling them out in the sampled navigation scenario. To ensure meaningful queries, colliding or self-intersecting trajectories filtered. We generate a dataset $D_{Q}$ of $N=500$ queries, from which we sample subsets for the participants to rate in the user study, either in VR or as 2D video playback.
@@ III-D Query Interfaces
To collect user preferences, we employ three types of query interfaces: an immersive virtual reality (VR) setup and two 2D video-based KBM interfaces on a desktop computer. Both interfaces show the same navigation environment but differ in perspective and immersion. The VR interface provides an interactive and immersive experience, whereas the 2D video KBM interface offers a more conventional, screen-based alternative. Within the 2D interface, we present two perspectives: Top-down (2D-TD), as in [24], also known as a bird’s-eye view, and first-person view (2D-FPV), which more closely resembles a VR perspective. For a given query, the human position is defined, serving additionally as the observer position in VR and 2D-FPV.
@@ III-D1 Virtual Reality
The VR interface is based on PyBullet VR [30], ensuring native compatibility with the EnQuery training environments. We connect an HTC Vive Pro Eye setup as VR hardware. A transparent blue cylinder indicates the static human observer position on the floor. Additionally, a floating dialogue in front of the user conveys instructions and announces the upcoming trajectory with labels (A or B) for $2\text{}\mathrm{s}$. For preference selection, participants interact with a floating selection menu by pointing and clicking on one of two labeled boxes, A or B.
@@
Each video begins with a $2\text{}\mathrm{s}$ trajectory label (A or B), ensuring clear differentiation between the two options. While the 2D-FPV video is recorded with a $60\text{}\mathrm{\SIUnitSymbolDegree}$ vertical field of view that tracks the robot from the perspective of the human, the 2D-TD perspective is set to capture the entire navigation path. Additionally, in 2D-TD, the human is represented by a neutral wooden mannequin 3D model.
@@ III-E User Study
We conducted a user study to compile a dataset of participants’ navigation preferences and assess user experience across three preference interfaces (VR, 2D-TD, 2D-FPV). Data collection was divided into three distinct stages: S1: Collecting preferences through navigation queries for each interface modality, S2: Post-interface interaction questionnaires assessing user experience, and S3: A final ranking survey comparing the three interfaces. Before testing, all participants received detailed study information, provided written consent, and completed a demographic questionnaire. The study was structured into three blocks in randomized order, each corresponding to one of the interfaces as an experimental condition. Each block presented the same 25 preference queries in random order (S1), initially sampled for each participant from the query dataset $D_{Q}$. By presenting identical queries across different interfaces, we could later investigate the impact of the interface modality on participants’ navigation preferences. After each interface block, participants completed a questionnaire (S2) regarding their experience with the interface and the queries, as shown in Fig. 2. Upon completion of all three blocks, the final ranking survey (S3), based on the Technology Acceptance Model (TAM) [32], asked participants to rank the interfaces by perceived usefulness, intention to use, and ease of use, as shown in I. All collected data were anonymized using a coding table for participant IDs. Each session lasted approximately one hour.
@@ IV Experimental Evaluation
We propose the following hypotheses: H1: The user experience differs between the interface modalities. H2: The user preferences for robot navigation differ between the interface modalities. H3: A preference discrepancy between interfaces reflects in the navigation behavior of interface-specific preference-aligned policies.
@@ IV-A Interface Questionnaire
Figure 2: Survey results (S2) comparing user experiences across three interface conditions: virtual reality (VR), 2D top-down (2D-TD), and 2D first-person view (2D-FPV). Participants rated their experience across multiple aspects after each block. Ratings were provided on a Likert scale (1-7), bars indicate score means, standard errors are indicated. Asterisks denote significance levels (* $p<.05$, ** $p<.01$, *** $p<.001$).
@@ IV-C Dataset Overview
The dataset is publicly available [^1] https://github.com/HumanoidsBonn/rlhf_prefnav_interface_study and contains 2,325 user preference queries for robot navigation, collected across three interface modalities: VR, 2D-TD, and 2D-FPV.
The dataset is structured by participant ID, interface modality, query, and preference. Each query stores both trajectories and the scene configuration (e.g., for replay in VR), the 2D videos, and RL episodes (state, action, next state, reward) for both trajectories, where states $\mathbf{s}_{t}=(\text{lidar},\text{goal},\text{human})$ capture lidar data, robot-centric goal and human position, while actions $\mathbf{a}_{t}=(v,\omega)$ represent velocity commands. Preferences are stored as A/B labels. In addition to the state-action pairs, we provide the robot trajectory as a 2D path, the static human pose, and the obstacle poses in world coordinates.
As we later show, our dataset enables the distillation of preference models for policy alignment.
@@ IV-D2 Disagreement Analysis
In cases where participants exhibited different preferences for the same queries in VR and on 2D interfaces, participants preferred shorter and more straightforward driving trajectories (measured by the area under the path) in 2D compared to VR. Additionally, the robot may traverse closer to the human, compare (e). Similarly, query disagreements point to a preference for more straightforward driving styles when transitioning from the 2D-FPV to the 2D-TD interface, see (f). Consequently, we find empirical support for H2.
@@ IV-E Policy Alignment
Steps 49.1 49.5 55.9 50.0  bcd
Path Length [m] 4.9 4.9 5.4 4.9  bcd
Time [s] 9.6 9.7 11.0 9.8  bcd
AA [rad] 3.2 3.1 4.1 2.8  bcd
AUP [m²] 2.2 1.8 3.4 2.1 abcd
$\min(d_{h})$ [m] 1.33 1.20 1.28 1.25 a c
Speed $\bar{v}$ [m/s] 0.51 0.50 0.49 0.51  bcd
TABLE III: Quantitative metrics for our preference-aligned $\pi_{\text{VR}}$, $\pi_{\text{\mbox{2D-TD}}}$, and $\pi_{\text{\mbox{2D-FPV}}}$ and a non-aligned baseline $\pi_{\text{BL}}$, averaged over 1,000 trajectories in randomly sampled scene configurations. Significance pair abbreviations ($p<.05$) for unequal mean (Welch t-test, Bonf. corr.) are a: VR vs. 2D-TD, b: VR vs. 2D-FPV, c: 2D-TD vs. 2D-FPV, d: 2D-FPV vs. BL.
To examine how preferences collected from different interfaces impact policy alignment and navigation behavior, we employ a preference-based reinforcement learning (PbRL) approach. This method trains reward functions from the collected preference dataset, enabling the subsequent alignment of human-aware RL navigation policies for each query interface. Following [10, 9], we learn a parametric reward function $\hat{r}_{\psi}$ from human preferences. For policy optimization, we implement a PbRL algorithm following [10], using TD3 [29] as the base RL algorithm. Splitting up our preference dataset by interface, we train three individual reward functions modeled as an MLP with $[256,256,256]$ hidden units based on the queries collected from the participants. The policies for each condition were then trained for 500k time steps by weighting the original navigation task reward and their respective preference reward model as $r=\lambda\hat{r}_{\psi}+(1-\lambda)r_{\text{core}}$, with $\lambda=0.2$. The value of $\lambda$ was empirically determined through a grid search to ensure optimal training stability. Note that the preference reward ranges from -1 to +1, while the original task reward from of -10 to +20, with these extrema driven by the sparse terms of $r_{\text{core}}$. In addition to the trained policies $\pi_{\text{VR}}$, $\pi_{\text{\mbox{2D-TD}}}$, and $\pi_{\text{\mbox{2D-FPV}}}$, we also include a non-aligned baseline policy $\pi_{\text{BL}}$, trained on the same task using $r_{\text{core}}$ without preference-based rewards, for comparison.
@@
The quantitative results in Table III confirm that preference-aligned policies improve human-aware navigation, with $\pi_{\text{VR}}$ achieving the best balance between efficiency and safety by maintaining the highest human clearance (1.33 m) while ensuring low travel times. $\pi_{\text{\mbox{2D-FPV}}}$ prioritizes safety, exhibiting the lowest collision rate (4.9%) but at the cost of longer paths (5.4 m) and higher angular accumulation (4.1 rad). $\pi_{\text{\mbox{2D-TD}}}$ exhibits the most straight-forward navigation, reflected in the lowest area under path (AUP). These findings, supported by the qualitative analysis, highlight that $\pi_{\text{VR}}$ offers the most balanced navigation performance, while $\pi_{\text{\mbox{2D-FPV}}}$ navigates overly conservatively, and $\pi_{\text{BL}}$ is less cautious. Statistical significance on the pairwise differences between policies are marked in Table III in column “sign. pairs” with indicators abcd based on Welch’s $t$-tests with Bonferroni correction. In summary, we find support for H3.
@@ V Conclusion
Our study systematically examined how interface modality affects human preference collection for robot navigation and the resulting policy alignment. The results confirm that the choice of interface modality significantly impacts how users express preferences, perceive the interaction, and ultimately shape robot navigation behavior. The VR interface provided a more immersive and intuitive experience, leading to greater confidence and ease in preference expression. This aligns with prior findings on the benefits of immersive environments for user engagement [35]. However, the study also revealed that preferences were not entirely consistent between interfaces, with participants favoring shorter, more direct paths in 2D interfaces while exhibiting greater tolerance for curved trajectories with increased human clearance in VR. This suggests that the visualization and spatial representation of the robot’s movement influence user preferences, highlighting the importance of maintaining interface consistency during preference collection. The navigation policies trained on interface-specific preferences demonstrated noticeable differences in robot behavior. These findings highlight the necessity of considering interface effects in preference-based reinforcement learning, as user preference shifts due to modality changes can directly impact policy training outcomes. To support further research on these effects and the collected user preferences, our preference dataset is available to the research community.